# Almost mathematics of pointed symmetric monoidal model categories by Smith ideal theory

### v4

@@ 2.2. Homotopically idempotent commutative Smith ideals
    Write $\tilde{j}:I\otimes I\to I\otimes V\to V$ and $\tilde{I}=I\otimes I$. Then $\tilde{I}$ is a cofibrant object and $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ has a homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the associativity:
@@
    The tensor product $\tilde{I}\otimes\tilde{I}$ is a group like object. That is, the unit morphism $\tilde{I}\otimes\tilde{I}\to\Omega\Sigma(\tilde{I}\otimes\tilde{I})$ is a weak equivalence, where $\Omega$ is the pointed loop functor and $\Sigma$ the suspension functor on the pointed model category $\mathcal{M}$.
@@
Let $\mathcal{M}$ be a symmetric monoidal model category with a monoidal unit $V$, and $j:\tilde{I}\to V$ a homotopically idempotent commutative Smith ideal of $V$. Then the product $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ is a weak equivalence.
  • proof.
@@
Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal of the monoidal unit $V$. A morphism $f:M\to N$ in $\mathcal{M}$ is an almost weak equivalence if the induced morphism $\tilde{I}\otimes Q(f):\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(N)$ is a weak equivalence, and $\mathcal{M}^{\rm al}$ denotes the Bousfield localization of $\mathcal{M}$ by the class of almost weak equivalences.
@@
An object $M$ of $\mathcal{M}$ is almost local if the induced morphism $f^{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{2},\,M)\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{1},\,M)$ is an isomorphism. Dually, $M$ is almost colocal if, for any homotopically almost weak equivalence $f:N_{1}\to N_{2}$, the induced map $f_{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{1})\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{2})$ is an isomorphism.
@@
$\mathrm{Map}_{\mathcal{M}}(\tilde{j},\,E):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,E)\to\mathrm{Map}_{\mathcal{M}}(V,\,E)\to E$
@@
In Quillen’s note [Qui96, Definition 5.2], $I$-closed modules said to be closed. However, this paper considers closed monoidal model categories, so we use the terminology $I$-closed objects for given a homotopically idempotent Smith ideal $I$.
@@
    For any object $M$, for any object $M$, $\tilde{j}:\tilde{I}\to V$ induces an almost weak equivalence $\tilde{j}:\tilde{I}\otimes M\to V\otimes M\to M$.
@@
  • proof.
    To prove (1), consider $Q(\tilde{j}):\tilde{I}\to Q(V)\to V$ as a factorization, where the first morphism is a cofibration and the second is a trivial fibration. Then, by the unit condition of symmetric monoidal model categories, $\tilde{I}\otimes Q(V)\otimes Q(M)\to\tilde{I}\otimes V\otimes Q(M)$ is a weak equivalence by definition. Additionally, since $\tilde{I}\otimes\tilde{I}\to\tilde{I}\otimes Q(V)$ is a weak equivalence of cofibrant objects, and $-\otimes Q(M)$ is a left Quillen functor, by Ken Brown’s lemma [Hov99, p.6 Lemma 1.1.12], we have that $\tilde{I}\otimes\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(V)\otimes Q(M)$ is also a weak equivalence. Consequently, $\tilde{I}\otimes\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(M)$ is a weak equivalence.
@@
Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, $j:I\to V$ be a homotopically idempotent commutative Smith ideal and $\mathcal{M}^{\rm al}$ denoting the Bousfield localization by the collection of almost weak equivalences. Then, the internal-hom functor
@@
Furthermore, let $\mathcal{M}^{\rm fi}$ denote the homotopically essential image of $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$. Then, the induced adjunction
$\tilde{I}\otimes Q:\mathcal{M}^{\rm al}\to\mathcal{M}^{\rm fi}:\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$
@@
  • proof.
Let $L^{\rm al}:\mathcal{M}\to\mathcal{M}^{\rm al}$ be another Bousfield localization by almost weak equivalences. By the universality of $L^{\rm al}$, the natural transformation $\mathrm{Id}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ homotopically factors as
@@
By (2) in Proposition 2.10, the composition $\psi(L^{\rm al})\circ\varphi(L^{\rm al}):L^{\rm al}\to L^{\rm al}\circ L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))$ is an equivalence. Therefore, the induced transformation $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\varphi):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))$ is identified with a homotopy section of $\psi:L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$. Further, since the functor $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is homotopically idempotent, the factorization
$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\psi)\circ\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\varphi):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\circ\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$
implies that $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is a homotopy retraction of $L^{\rm al}$. Hence, the natural transformation $\psi:L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$ is a weak equivalence.
We show that the functor $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$ induced a left Quillen equivalence $\tilde{I}\otimes Q:\mathcal{M}^{\rm al}\to\mathcal{M}^{\rm fi}$. Since the functor $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$ sends almost weak equivalences to weak equivalences, it factors through $\mathcal{M}^{\rm al}$. By (1) in Proposition 2.10 and the equivalence $\psi:L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$, the unit $\mathrm{Id}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\circ(\tilde{I}\otimes Q)$ induces an equivalence on the homotopy category of $\mathcal{M}^{\rm al}$, implying that it is a left Quillen equivalence. $\Box$
@@
First, we consider the module category over a unital ring. Let $V$ be a commutative unital ring with an idempotent ideal $\mathfrak{m}$. Assume that $\tilde{\mathfrak{m}}$ is a flat $V$-module. In Smith ideal theory language, a $V$-homomorphism $\tilde{j}:\tilde{\mathfrak{m}}\to V$ is regarded a commutative Smith ideal of $V$. For any $V$-algebra $A$, the almost unitalization $A_{!!}$ can be represented by the pushout product
@@
which is an almost isomorphism. In Gabber–Ramero’s text [GR03], $A_{!!}$ is defined by to be the source of $\tilde{j}\Box(V\to A)$ without Smith ideal theory. On the view point of Smith ideal theory, we can easily check that the induced ring homomorphism $A_{!!}\to A$ is an almost isomorphism. Indeed, by the associativity of pushout products,
$(\tilde{\mathfrak{m}}\otimes_{V}{A}_{!!}\to\tilde{\mathfrak{m}}\otimes_{V}A)\simeq L_{1}(\tilde{\mathfrak{m}})\Box(\tilde{j}\Box(V\to A))\to(L_{1}(\tilde{\mathfrak{m}})\Box\tilde{j})\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}}\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}\otimes_{V}A}$
@@ 3.1. Almost unitalization functors of commutative algebra objects of symmetric monoidal model categories
In general the adjunction $\mathrm{Sym}:\mathcal{M}\rightleftarrows\mathrm{CAlg}(\mathcal{M}):U$ does not induces a model structure on $\mathrm{CAlg}(\mathcal{M})$ but it has the semi-model structure defined in [Spi01, Definition 2.3] or [Whi22, Definition 3.9]. Semi-model structure only needs that trivial cofibrations with cofibrant domains have the left lifting property with respect to all fibrations, and only maps with cofibrant domains factor into a trivial cofibration followed by a fibration. If all objects of a semi-model category are cofibrant, it is just a model category (See [Whi22, Definition 3.9]). Left and right Quillen functors between semi-model categories are defined as the same as ones of model categories.
For any class of morphisms $\mathcal{I}$, let $\mathrm{Cell}(\mathcal{I})$ denote the class of morphisms that is a transfinite composition of pushouts of the element of $\mathcal{I}$, and an object of $\mathrm{Cell}(\mathcal{I})$ is said to be a relative $\mathcal{I}$-cell complex.
Let $\mathcal{M}$ is cofibrantly generated with a set of generating cofibrations $\mathcal{I}$ and a set of generating trivial cofibrations $\mathcal{J}$. By White [Whi22, Corollary 3.8], if the domains of $\mathcal{I}$ (resp. $\mathcal{J}$) is small relative to $\mathrm{Cell}(\mathcal{I}\otimes\mathcal{M})$ (resp. $\mathrm{Cell}(\mathcal{J}\otimes\mathcal{M})$), then the category of commutative $R$-algebras has a semi-model structure inherited by $\mathcal{M}$ for any commutative monoid object $R$.
@@
Definition 3.1 ([SS03] Definition 3.6).
Let $F:\mathcal{M}\rightleftarrows\mathcal{N}:U$ be a Quillen adjunction of monoidal model categories. Then we say that $(F,\,U)$ is lax Quillen monoidal if it satisfies the following properties:
@@
  • proof.
    Since the class almost weak equivalences between cofibrant objects is closed under tensor product, the localization $L^{\rm al}:\mathcal{M}\to\mathcal{M}^{\rm al}$ is lax monoidal. Furthermore, given any pair of cofibrant objects $M$ and $N$, the product $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ induces a weak equivalence
@@
$\tilde{I}\otimes Q:\mathrm{CAlg}(\mathcal{M}^{\rm al})\rightleftarrows\mathrm{CAlg}(\mathcal{M}^{\rm fi}):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)$
@@
We verify that, for any commutative monoid object $A$, the internal-hom object $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)$ has a commutative monoid structure on $\mathcal{M}$ inherited by one of $A$ and the homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the coherent conditions in Definition 2.2. Let $\mathrm{ev}_{\tilde{I}}:\tilde{I}\otimes\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\to A$ denote the evaluation morphism induced by the identity on $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)$. Then the composition
@@
satisfying the associativity and unitality due to Condition (3) in Definition 2.2.
In the case that the monoidal structure on $\mathcal{M}$ is closed, both the symmetric monoidal structures on the allow category $\mathcal{M}$ are closed by Hovey [Hov14, Theorem 1.2]. For any couple of morphisms $(f:X_{0}\to X_{1},\,g:Y_{0}\to Y_{1})$ in $\mathcal{M}$, the internal-hom objects are determined as follows: The internal-hom object $\mathrm{Map}_{\mathcal{M}}^{\otimes}(f,\,g)$ of the diagonal product monoidal structure is defined by
@@
Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category with a monoidal unit object $V$ and homotopically idempotent commutative Smith ideal $j:I\to V$. Assume that $\mathcal{M}$ is left proper and combinatorial. Then the Quillen adjunction
$\tilde{j}\Box(-):\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}))\rightleftarrows\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M})):\mathrm{Map}^{\Box}(\tilde{j},\,-)$
@@
$\tilde{j}\Box(-):\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al}))\rightleftarrows\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al})):\mathrm{Map}^{\Box}(\tilde{j},\,-),$
@@
By Proposition 3.5, $j\Box\mathrm{Id}\to\mathrm{id}_{Ev_{1}(-)}=L_{0}(\mathrm{Ev}_{1}((-)))$ is a projective equivalence in $\mathrm{CAlg}(\mathcal{M}^{\rm al})$. Therefore, their Quillen right adjoints $\mathrm{Map}^{\Box}(\tilde{j},\,-)$ and $U_{1}\circ\mathrm{Ev}_{0}$ are weakly equivalent on $\mathrm{Ar}(\mathrm{CAlg}(\mathcal{M}^{\rm al}))$. $\Box$
We obtain an analogue of Gabber–Ramero [GR03, p.22, Proposition 2.2.29] as avv corollary of Theorem 3.6.
@@
  • proof.
    Since all of $L_{1}$, $\tilde{j}\Box(-)$, and $\mathrm{Ev}_{0}$ are left Quillen functors, the composition $(-)_{!!}$ is also a left Quillen functor, and its right adjoint is the composition of their right adjoint functors

### v5

@@ 2.2. Homotopically idempotent commutative Smith ideals
    Write $\tilde{I}=I\otimes I$. Then $\tilde{I}$ is cofibrant and $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ has a homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the associativity:
@@
[…]
@@
Let $\mathcal{M}$ be a symmetric monoidal model category with a monoidal unit $V$, and $j:I\to V$ a homotopically idempotent commutative Smith ideal of $V$. Then the product $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ is a weak equivalence.
@@
Example 2.6 ([6, 12] Torsion and derived completion).
Let $A$ be a commutative ring and $\mathfrak{a}\subset A$ a finitely generated weakly proregular ideal. Consider the injective model category $\operatorname{Ch}(A)$ and let $K_{\infty}(\mathfrak{a})$ denote the infinite dual Koszul complex. Then, on the derived category $D(A)$, the derived torsion functor $\mathbb{R}\Gamma_{\mathfrak{a}}(-)$ is equivalent to the tensor functor $K_{\infty}(\mathfrak{a})\otimes_{A}^{\mathbb{L}}(-)$, and the derived completion functor $\mathbb{L}\Lambda_{\mathfrak{a}}(-)$ is equivalent to the internal-hom functor $\mathrm{Map}_{A}(K_{\infty}(\mathfrak{a}),\,-)$. Note that $K_{\infty}(\mathfrak{a})$ is homotopy idempotent: $K_{\infty}(\mathfrak{a})\otimes_{A}^{\mathbb{L}}K_{\infty}(\mathfrak{a})\simeq K_{\infty}(\mathfrak{a})$. Therefore, the adjunction
$K_{\infty}(\mathfrak{a})\otimes_{A}^{\mathbb{L}}(-):D(A)\rightleftarrows D(A):\mathrm{Map}_{A}(K_{\infty}(\mathfrak{a}),\,-)$
gives a bilocalization of $D(A)$, where derived torsion is realized as the tensor side and derived completion as the internal-hom side. In this sense, the derived completion functor is an instance of the internal-hom side of our general tensor–internal-hom bilocalization. This viewpoint is compatible with the equivalence between firm and closed modules over an idempotent ideal: derived $\mathfrak{a}$-torsion complexes are the fixed objects of $K_{\infty}(\mathfrak{a})\otimes_{A}^{\mathbb{L}}(-)$, while derived $\mathfrak{a}$-complete complexes are the fixed objects of $\mathrm{Map}_{A}(K_{\infty}(\mathfrak{a}),\,-)$. The adjunction restricts to the Greenlees–May duality between the corresponding full subcategories [12, 13]. Hence the commutative Smith ideal $j:I\to A$ obtained from a cofibrant replacement of $K_{\infty}(\mathfrak{a})\to A$ gives a fundamental example of a homotopically idempotent commutative Smith ideal governing this duality.
@@
Let $j:I\to V$ be a homotopically idempotent commutative Smith ideal of the monoidal unit $V$. A morphism $f:M\to N$ in $\mathcal{M}$ is called an almost weak equivalence if the induced morphism $\tilde{I}\otimes Q(f):\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(N)$ is a weak equivalence, and $\mathcal{M}^{\rm al}$ denotes the Bousfield localization of $\mathcal{M}$ with respect to the class of almost weak equivalences.
@@
An object $M$ of $\mathcal{M}$ is almost local if, for any almost weak equivalence $f:N_{1}\to N_{2}$, the induced morphism
$f^{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{2},\,M)\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(N_{1},\,M)$
is an isomorphism. Dually, $M$ is almost colocal if, for any almost weak equivalence $f:N_{1}\to N_{2}$, the induced map
$f_{*}:\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{1})\to\mathrm{Hom}_{\mathrm{Ho}\mathcal{M}}(M,\,N_{2})$
is an isomorphism.
@@
$\mathrm{Map}_{\mathcal{M}}(\tilde{j},\,E):\mathrm{Map}_{\mathcal{M}}(V,\,E)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,E)$
@@
In Quillen’s note [14, Definition 5.2], $I$-closed modules are said to be closed. However, this paper considers closed monoidal model categories, so we use the terminology $I$-closed objects for a given homotopically idempotent Smith ideal $I$.
@@
    For any object $M$, $\tilde{j}:\tilde{I}\to V$ induces an almost weak equivalence $\tilde{j}:\tilde{I}\otimes M\to V\otimes M\to M$.
@@
To prove (1), consider $Q(\tilde{j}):\tilde{I}\to Q(V)\to V$ as a factorization, where the first morphism is a cofibration and the second is a trivial fibration. Then, by the unit condition of symmetric monoidal model categories, $\tilde{I}\otimes Q(V)\otimes Q(M)\to\tilde{I}\otimes V\otimes Q(M)$ is a weak equivalence by definition. Additionally, since $\tilde{I}\otimes\tilde{I}\to\tilde{I}\otimes Q(V)$ is a weak equivalence of cofibrant objects, and $\otimes Q(M)$ is a left Quillen functor, by Ken Brown’s lemma [8, p.6 Lemma 1.1.12], we have that $\tilde{I}\otimes\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(V)\otimes Q(M)$ is also a weak equivalence. Consequently, $\tilde{I}\otimes\tilde{I}\otimes Q(M)\to\tilde{I}\otimes Q(M)$ is a weak equivalence.
@@
Let $\mathcal{M}$ be a pointed symmetric monoidal model category with a monoidal unit object $V$, let $j:I\to V$ be a homotopically idempotent commutative Smith ideal, and let $\mathcal{M}^{\rm al}$ denote the Bousfield localization with respect to the collection of almost weak equivalences. Then the internal-hom functor
@@
Furthermore, let $\mathcal{M}^{\rm fi}$ denote the homotopically essential image of $\tilde{I}\otimes Q:\mathcal{M}\to\mathcal{M}$. Then the induced adjunction
[…]
@@
[…]
By (2) in Proposition 2.11, the composition $\psi(L^{\rm al})\circ\varphi(L^{\rm al}):L^{\rm al}\to L^{\rm al}\circ L^{\rm al}\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))$ is an equivalence. Therefore, the induced transformation
$\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,\varphi):\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,-)\to\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,L^{\rm al}(-))$
[…]
@@
First, we consider the module category over a unital ring. Let $V$ be a commutative unital ring with an idempotent ideal $\mathfrak{m}$. Assume that $\tilde{\mathfrak{m}}$ is a flat $V$-module. In the language of Smith ideal theory, a $V$-homomorphism $\tilde{j}:\tilde{\mathfrak{m}}\to V$ is regarded as a commutative Smith ideal of $V$. For any $V$-algebra AA, the almost unitalization $A_{!!}$ can be represented by the pushout product
@@
which is an almost isomorphism. In Gabber–Ramero’s text [5], $A_{!!}$ is defined as the source of $\tilde{j}\Box(V\to A)$ without using Smith ideal theory. From the viewpoint of Smith ideal theory, one can easily check that the induced ring homomorphism $A_{!!}\to A$ is an almost isomorphism. Indeed, by the associativity of pushout products,
$(\tilde{\mathfrak{m}}\otimes_{V}{A}_{!!}\to\tilde{\mathfrak{m}}\otimes_{V}A)\simeq L_{1}(\tilde{\mathfrak{m}})\Box(\tilde{j}\Box(V\to A))\\ \to(L_{1}(\tilde{\mathfrak{m}})\Box\tilde{j})\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}}\Box(V\to A)\simeq\mathrm{id}_{\tilde{\mathfrak{m}}\otimes_{V}A}$
@@ 3.1. Almost unitalization functors of commutative algebra objects of symmetric monoidal model categories
In general the adjunction $\mathrm{Sym}:\mathcal{M}\rightleftarrows\mathrm{CAlg}(\mathcal{M}):U$ does not induces a model structure on $\mathrm{CAlg}(\mathcal{M})$ but it has the semi-model structure defined in [17, Definition 2.3] or [19, Definition 3.9]. A semi-model structure only requires that trivial cofibrations with cofibrant domains have the left lifting property with respect to all fibrations, and only maps with cofibrant domains factor into a trivial cofibration followed by a fibration. If all objects of a semi-model category are cofibrant, it is just a model category (see [19, Definition 3.9]). Left and right Quillen functors between semi-model categories are defined in the same way as those of model categories.
For any class of morphisms $\mathcal{I}$, let $\mathrm{Cell}(\mathcal{I})$ denote the class of morphisms that are transfinite compositions of pushouts of elements of $\mathcal{I}$, and a morphism in $\mathrm{Cell}(\mathcal{I})$ is said to be a relative $\mathcal{I}$-cell complex.
Let $\mathcal{M}$ be cofibrantly generated with a set of generating cofibrations $\mathcal{I}$ and a set of generating trivial cofibrations $\mathcal{J}$. By White [19, Corollary 3.8], if the domains of $\mathcal{I}$ (resp. $\mathcal{J}$) are small relative to $\mathrm{Cell}(\mathcal{I}\otimes\mathcal{M})$ (resp. $\mathrm{Cell}(\mathcal{J}\otimes\mathcal{M})$), then the category of commutative $R$-algebras has a semi-model structure inherited from $\mathcal{M}$ for any commutative monoid object $R$.
@@
Definition 3.1 ([16] Definition 3.6).
Let $F:\mathcal{M}\rightleftarrows\mathcal{N}:U$ be a Quillen adjunction of monoidal model categories. Then we say that (F,U)(F,\,U) is lax Quillen monoidal if it satisfies the following properties:
@@
Since the class of almost weak equivalences between cofibrant objects is closed under tensor product, the localization $L^{\rm al}:\mathcal{M}\to\mathcal{M}^{\rm al}$ is lax monoidal. Furthermore, given any pair of cofibrant objects $M$ and $N$, the product $\mu_{\tilde{I}}:\tilde{I}\otimes\tilde{I}\to\tilde{I}$ induces a weak equivalence
@@
[…]
@@
We verify that, for any commutative monoid object $A$, the internal-hom object $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)$ has a commutative monoid structure in $\mathcal{M}$ inherited from that of $A$ and the homotopy section $\Delta:\tilde{I}\to\tilde{I}\otimes\tilde{I}$ satisfying the coherent conditions in Definition 2.2. Let $\mathrm{ev}_{\tilde{I}}:\tilde{I}\otimes\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)\to A$ denote the evaluation morphism induced by the identity on $\mathrm{Map}_{\mathcal{M}}(\tilde{I},\,A)$. Then the composition
@@
satisfying associativity and unitality by Condition (3) in Definition 2.2.
In the case that the monoidal structure on $\mathcal{M}$ is closed, both symmetric monoidal structures on the arrow category $\mathrm{Ar}(\mathcal{M})$ are closed by Hovey [9, Theorem 1.2]. For any pair of morphisms $(f:X_{0}\to X_{1},\,g:Y_{0}\to Y_{1})$ in $\mathcal{M}$, the internal-hom objects are determined as follows: The internal-hom object $\mathrm{Map}_{\mathcal{M}}^{\otimes}(f,\,g)$ of the diagonal product monoidal structure is defined by
@@
Let $\mathcal{M}$ be a pointed closed symmetric monoidal model category, and $V$ a monoidal unit of $\mathcal{M}$ with a homotopically idempotent commutative Smith ideal $j:I\to V$. Assume that $\mathcal{M}$ is left proper and combinatorial. Then the Quillen adjunction
[…]
@@
[…]
@@
[…]
We obtain an analogue of Gabber–Ramero [5, p.22, Proposition 2.2.29] as a corollary of Theorem 3.6.
@@
[…]